# Photoacoustic Iterative Optimization Algorithm with Shape Prior Regularization

### v2

@@ 2.2.2 Optimizer Selection and Regularized Iteration Process
2.2.2.1 Optimizer Selection
We selected the Adam optimizer (Adaptive Moment Estimation)[21] as our optimization algorithm. Adam is a first-order gradient-based optimization algorithm that combines the advantages of Momentum and RMSProp, making the optimization process more stable and efficient in complex reconstruction tasks. Specifically, the momentum mechanism reduces oscillations and gradient noise by applying an exponentially weighted average to the gradients, while RMSProp’s adaptive learning rate ensures proper adjustments of the learning rates across different variable dimensions. Additionally, Adam incorporates bias correction, which rectifies the biases during early iterations, thereby improving stability. Overall, with its fast convergence and adaptive features, the Adam optimizer meets the requirements of our task.
2.2.2.2 Regularized Iteration Process
@@
3:          $P_{N}$: Probability matrix     $\triangleright$ $[0,1]$
@@
12:               $gradient_{prior\_reg}\leftarrow 2\times(1-P_{N})\times R_{op}$
@@
3 Result
In this part, we use 3D simulation data and in vivo animal experiment data to verify the performance of RI-SPPM algorithm.
@@ 3.1 Simulation Data Results
We first used simulated hand vessels as the target tissue to demonstrate the high performance of the RI-SPPM algorithm. The simulated data is based on the 3D PA imaging results of human peripheral hand vessels from another work[22], which was used as the ground truth. At a grid point spacing of 0.2 mm, the total grid size of the imaging area is 128*320*256 (25.6 mm ×64.0 mm ×51.2 mm). In order to simulate 3D PA imaging, the radius of the hemispherical system is 4cm, and a gap with a height of 1cm is opened at the top as the laser light inlet, and the remaining positions are evenly distributed with 1024 array elements(Figure 3 is imaging setup). Then we obtained PA signals collected by these 1024 arrays through forward propagation of the K-Wave toolbox[23].
(a) Imaging setup
(b) Imaging setup’s another view
Figure 3: (a) shows the approximate position of the elements and fingers in the system. (b) further shows the position of the fingers and the direction of placement
@@
The above imaging system was selected for simulation to generate sparse photoacoustic signals in order to evaluate the performance of the RI-SPPM algorithm. The simulation is based on 3D photoacoustic imaging data of real hand vasculature as the "ground truth," using a high-resolution grid with a resolution of 0.2 mm (128×320×256) to capture the intricate structures of the vasculature. The system is designed as a hemispherical array with a radius of 4 cm, featuring a laser entry point at the top with a height of 1 cm to create a missing-view angle. The array is uniformly distributed with 1024 sensors to simulate a typical photoacoustic imaging device. This simulation approach closely approximates real-world scenarios, provides a controlled evaluation standard, and offers a reliable basis for validating the performance of the RI-SPPM algorithm in photoacoustic imaging.
3.1.1.1 Comparison of Results before and after Optimization
@@
In this example, we use the 3D PA result reconstructed by the UBP algorithm as the result optimized by the RI-SPPM algorithm. The distributions of the acoustic source are presented in Figure 4(a), while the reconstruction results are depicted in Figure 4(b-c). For each set of images, the upper and lower subimages correspond to the top-view maximum amplitude projection (MAP) and the front-view MAP of the 3D reconstruction, respectively. In Figure 4(b), the reconstruction results obtained via the UBP algorithm display significant artifacts on the right side of the index finger. These arise because the laser illumination originates from the right side of the index finger, leading to a lack of information from that particular angle and, consequently, severe artifacts. When compared to the ground truth, a substantial discrepancy is evident. As shown in Figure 4(c), the optimization performed by the RI-SPPM algorithm significantly reduces the artifacts in the UBP results, bringing them closer to the ground truth. The artifacts caused by incomplete angular coverage are largely corrected, and additional spurious artifacts in other regions are also mitigated, such as those observed in the middle of the finger’s top-view MAP and in the upper and lower areas of the finger’s front-view MAP. Furthermore, in the 3D view (Figure 4(e-f)), the improvements in the PA reconstruction results after applying the RI-SPPM algorithm are clearly observable.
(a) Ground truth
(b) UBP reconstruction
(c) RI-SPPM optimization
(d) UBP restruction
(e) RI-SPPM optimization
@@
(c) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RI-SPPM result. (d) Reconstruction result of UBP algorithm in 3D-view. (e) Reconstruction result after optimization of RI-SPPM algorithm in 3D-view.
Subsequently, we performed a comprehensive mean squared error (MSE) analysis on the reconstruction results before and after optimization (all three datasets are normalized). The MSE between the reconstruction using the UBP algorithm and the ground truth is $0.3398$, whereas the MSE after applying the RI-SPPM optimization is reduced to $0.0040$. To further demonstrate the effectiveness of the RI-SPPM optimization, and considering that the overall MSE metric may be influenced by the presence of artifacts, we selected a region rich in valid information for detailed analysis and verification. Specifically, we extracted data along two lines. The voxel intensity values along the lines for the three datasets are presented in Figure 5. In Figure 5(b), it is evident that, in regions containing finger information, the intensity curve of the RI-SPPM-optimized result aligns closely with the ground truth, demonstrating superior consistency compared to the UBP-only reconstruction. The correlations before and after using RI-SPPM are $0.65$ and $0.92$. In Figure 5 (c), it can be seen that in regions lacking finger data, the noise and artifact levels in the RI-SPPM-optimized results are significantly closer to zero, while the noise and artifact levels in the UBP-only reconstruction are considerably higher. The correlations before and after using RI-SPPM are $0.00$ and $0.87$.
(a) The location of the two lines used for data analysis
(b) With finger information
(c) No finger information
@@
In conclusion, in the simulation experiments, the reconstruction results obtained using the UBP algorithm and further optimized by the RI-SPPM algorithm demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
3.1.1.2 Parameters Selection and Computing Environment
@@
Table 1: 3D PA image reconstruction under 1024 hemispherical system
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor x 128. In UBP algorithm reconstruction, we use Taichi[24, 25] for GPU acceleration. For the reconstruction with PA signal size of 1024*4096 and region size of 128*320*256, the time consuming is about 0.9s. So the entire RI-SPPM optimization process took less than a minute and 30 seconds. Of course, we have skills in the choice of parameters, and the choice of x is generally about one-tenth of the total number of elements. We also found that when the initial value of lr is about 0.001, the iterative result can be quickly optimized. And the number of iterations is much lower than num_iters.
3.1.2 3D PA Image Reconstruction under 256 Spherical System
We then used a Gaussian sphere and three cylinders as target to demonstrate the high performance of the RI-SPPM algorithm (Figure 6(a)). At a grid point spacing of 0.1 mm, the total grid size of the imaging area is 256*256*256 (25.6 mm ×25.6 mm ×25.6 mm). In order to simulate 3D PA imaging, the radius of the spherical system is 6cm. Then we obtained PA signals collected by these 256 arrays through forward propagation of the K-Wave toolbox.
3.1.2.1 Comparison of Results before and after Optimization
From a three-dimensional perspective (Figure 6(b,c)), the artifacts significantly decrease after optimization by the RI-SPPM algorithm. The Gaussian sphere and tubular structures are displayed more clearly. Through calculations, the MSE between the UBP reconstructed result and the ground truth is $0.0126$, while the MSE of the result optimized by the RI-SPPM algorithm is $0.000146$. In conclusion, in the simulation experiments, the reconstruction results obtained using the UBP algorithm and further optimized by the RI-SPPM algorithm demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
(a) Ground truth
(b) UBP reconstruction
(c) RI-SPPM optimization
Figure 6: Comparison of 3D photoacoustic reconstruction results under 256 spherical system. (a) Ground truth. (b) Reconstruction result of UBP algorithm in 3D-view. (c) Reconstruction result after optimization of RI-SPPM algorithm in 3D-view.
To better show the contrast, in Figure 7 we use a jet map to display the maximum amplitude projection in the direction of the red arrow in Figure 6(a).
(a) Ground truth
(b) UBP reconstruction
(c) RI-SPPM optimization
Figure 7: Comparison of 3D photoacoustic reconstruction results under 256 spherical system. (a) The maximum amplitude projection of ground truth. (b) The maximum amplitude projection of the UBP result. (c) The maximum amplitude projection of the RI-SPPM result.
3.1.2.2 Parameters Selection and Computing Environment
@@
$num\_iters$ 500
Table 2: 3D PA image reconstruction under 256 spherical system
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor x 128. In UBP algorithm reconstruction, we use Taichi for GPU acceleration. For the reconstruction with PA signal size of 256*4096 and region size of 256*256*256, the time consuming is about 0.8s. So, the entire RI-SPPM optimization process took less than a minute and 30 seconds. Of course, we have skills in the choice of parameters, and the choice of x is generally about one-tenth of the total number of elements. We also found that when the initial value of lr is about 0.001, the iterative result can be quickly optimized. The number of iterations is much lower than num_iters.
@@ 3.2 In Vivo Human Experiment Results
Then in vivo human study data, including human arm was acquired by Shuang Li’s work[22] using synthetic matrix array. The system employed a nonfocusing linear array (customized by Imasonics, France) to receive PA signals. The linear array has 256 elements with a pitch of 0.5 mm and a kerf of 0.1 mm, i.e., a total length of 12.8 cm. The center frequency of the ultrasonic array is 3.5 MHz with over 80% bandwidth. PA signal is amplified 1500 times via self-built two-stage amplifier, and then received by a data acquisition system (Marsonics DAQ, Tianjin Langyuan Technology Co., Ltd. China) at 40 MHz sampling rate. The system is shown in Figure 7.
(a) 3D PA imaging system
(b) Schematic diagram of fiber optic illumination and one-dimensional array
Figure 8: (a) shows the 3D PA imaging system. (b) shows details of the core of the system [22]
In the original work, a linear array consisting of 256 elements was moved 2969 times, with a step size of 0.1 mm for each movement, resulting in a large-scale synthetic matrix of 256 × 2969 elements. To create a sparse-view setup for better evaluation of the RI-SPPM algorithm’s performance, we extracted 120 rows at equal intervals (approximately 25% of the original data), which corresponds to a step size of 2.5 mm for each scan of the linear array.
@@ 3.2.2 Comparison of Results before and after Optimization
In this experiment, we utilized the RI-SPPM algorithm to optimize the reconstruction results obtained from the UBP algorithm. The detailed comparison of results before and after optimization is presented in Figure 8.
We compared the hand imaging results before and after the RI-SPPM optimization of the UBP reconstruction. From the top-view perspective, the optimization significantly reduced noise, and both thick and thin blood vessels were displayed more clearly. From the front-view perspective, in addition to making the vascular information more distinct, the lateral noise at the bottom was also effectively suppressed. These findings demonstrate that the reconstruction quality after RI-SPPM optimization is exceptionally high.
In conclusion, in vivo human experiment, the reconstruction results obtained using the UBP algorithm and further optimized by the RI-SPPM algorithm also demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
(a) UBP reconstruction
(b) RI-SPPM optimization
@@
(b) Hand’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RI-SPPM results.
@@ 3.2.3 Parameters Selection and Computing Environment
Table 3: 3D PA image reconstruction under synthetic matrix array
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor x 128. In UBP algorithm reconstruction, we use Taichi for GPU acceleration. For the reconstruction with PA signal size of 30720*2048 and region size of 625*600*250, the time consuming is about 10.2s. For the reconstruction with PA signal size of 4000*2048 and region size of 625*600*250, the time consuming is about 3.5s. So the entire RI-SPPM optimization process took less than 7 minutes.
@@ 4 Discussion and Future
In this study, we focus on demonstrating the optimization effectiveness of the RI-SPPM algorithm applied to the reconstruction results of the UBP algorithm in the context of 3D imaging systems. Through a series of specific experiments, the RI-SPPM algorithm has shown exceptional performance in both simulated studies and live small animal experiments. Notably, from a theoretical perspective, the RI-SPPM algorithm is not constrained to a specific imaging dimension, making it applicable to both 3D and 2D imaging scenarios. Moreover, the core concept of the RI-SPPM algorithm lies in optimizing existing imaging results rather than directly reconstructing images within the signal domain. This unique characteristic grants the algorithm broad applicability, enabling effective suppression of noise and artifacts caused by transducer-related factors (e.g., transducer distribution and other system properties). Importantly, the RI-SPPM optimization is not tied to a specific reconstruction algorithm; other algorithms can also benefit from its enhancement capabilities.
Furthermore, the potential applications of the RI-SPPM algorithm extend beyond photoacoustic computed tomography. Theoretical analysis suggests that the algorithm could be generalized to other imaging fields, such as ultrasound imaging, computed tomography (CT), and various other medical imaging or industrial inspection scenarios. In these fields, the sampling and imaging processes of various devices may also introduce artifacts and noise that vary with the properties of the transducers. In such cases, the RI-SPPM algorithm is expected to offer significant assistance.
Going forward, we plan to expand our research in two key directions. Firstly, we aim to further investigate the optimization capabilities of the RI-SPPM algorithm across a broader spectrum of existing reconstruction algorithms, exploring its adaptability under diverse imaging conditions, modalities, and algorithmic outputs. Secondly, we intend to extend the application of the RI-SPPM algorithm to other areas, such as ultrasound imaging, CT, and potentially other advanced medical imaging techniques like magnetic resonance imaging (MRI) and positron emission tomography (PET). These efforts would further validate its extensive practical value and generalizability.
In the long term, we aspire to make the RI-SPPM algorithm an essential tool for a wide range of imaging systems. Through continued optimization and extension studies, this algorithm has the potential to provide comprehensive solutions for improving imaging quality, resolution, and reliability across diverse application scenarios.

### v4

@@ 2.2.2 Optimizer Selection and Regularized Iteration Process
Optimizer Selection
We selected the Adam optimizer (Adaptive Moment Estimation) as our optimization algorithm. Adam is a first-order gradient-based optimization algorithm that combines the advantages of momentum and root mean square propagation (RMSProp)[20], making the optimization process more stable and efficient in complex reconstruction tasks. Specifically, the momentum mechanism reduces oscillations and gradient noise by applying an exponentially weighted average to the gradients, while RMSProp’s adaptive learning rate ensures proper adjustments of the learning rates across different variable dimensions. Additionally, Adam incorporates bias correction, which rectifies the biases during early iterations, thereby improving stability. Overall, with its fast convergence and adaptive features, the Adam optimizer meets the requirements of our task.
Regularized Iteration Process
@@
3:          $P_{norm}$: Probability matrix     $\triangleright$ $[0,1]$
@@
12:               $gradient_{prior\_reg}\leftarrow 2\times(1-P_{N})\odot R_{op}$
@@
3 Results
In this session, we use 3D simulation data and in vivo animal experiment data to verify the performance of RISP algorithm.
@@ 3.1 Simulation Data Results
We first used simulated hand vessels as the target tissue to demonstrate the high performance of the RISP algorithm. The simulated data is based on the 3D PA imaging results of human peripheral hand vessels from another work[21], which was used as the ground truth. At a grid point spacing of 0.2 mm, the total grid size of the imaging area is 128$\times$320$\times$256 (25.6 mm $\times$64.0 mm $\times$51.2 mm). In order to simulate 3D PA imaging, the radius of the hemispherical system is 4cm, and a gap with a height of 1cm is opened at the top as the laser light inlet, and the remaining positions are evenly distributed with 1024 array elements(Figure 1 is imaging setup). Then we obtained PA signals collected by these 1024 arrays through forward propagation of the K-Wave toolbox[22].
Figure 1: The position of the fingers and sensors
@@
The above imaging system was selected for simulation to generate sparse photoacoustic signals in order to evaluate the performance of the RISP algorithm. The simulation is based on 3D photoacoustic imaging data of real hand vasculature as the "ground truth," using a high-resolution grid with a resolution of 0.2 mm (128×320×256) to capture the intricate structures of the vasculature. The system is designed as a hemispherical array with a radius of 4 cm, featuring a laser entry point at the top with a height of 1 cm to create a missing-view angle. The array is uniformly distributed with 1024 sensors to simulate a typical photoacoustic imaging device. This simulation approach closely approximates real-world scenarios, provides a controlled evaluation standard, and offers a reliable basis for validating the performance of the RISP algorithm in photoacoustic imaging.
Comparison of Results before and after Optimization
@@
In this example, we use the 3D PA result reconstructed by the UBP algorithm as the result optimized by the RISP algorithm. The distributions of the acoustic source are presented in Figure 2(a), while the reconstruction results are depicted in Figure 2(b-c). For each set of images, the upper and lower subimages correspond to the top-view maximum amplitude projection (MAP) and the front-view MAP of the 3D reconstruction, respectively. In Figure 2(b), the reconstruction results obtained via the UBP algorithm display significant artifacts on the right side of the index finger. These arise because the laser illumination originates from the right side of the index finger, leading to a lack of information from that particular angle and, consequently, severe artifacts. When compared to the ground truth, a substantial discrepancy is evident. As shown in 2(c), the optimization performed by the RISP algorithm significantly reduces the artifacts in the UBP results, bringing them closer to the ground truth. The artifacts caused by incomplete angular coverage are largely corrected, and additional spurious artifacts in other regions are also mitigated, such as those observed in the middle of the finger’s top-view MAP and in the upper and lower areas of the finger’s front-view MAP. Furthermore, in the 3D view (Figure 2(e-f)), the improvements in the PA reconstruction results after applying the RISP algorithm are clearly observable.
@@
(c) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RISP result. (d) Reconstruction result of UBP algorithm in 3D-view. (e) Reconstruction result after optimization of RISP algorithm in 3D-view.
Specifically, we extracted data along two lines. The voxel intensity values along the lines for the three datasets are presented in Figure 3. In Figure 3(b), it is evident that, in regions containing finger information, the intensity curve of the RISP-optimized result aligns closely with the ground truth, demonstrating superior consistency compared to the UBP-only reconstruction. The correlations before and after using RISP are $0.65$ and $0.92$. In Figure 3(c), it can be seen that in regions lacking finger data, the noise and artifact levels in the RISP-optimized results are significantly closer to zero, while the noise and artifact levels in the UBP-only reconstruction are considerably higher. The correlations before and after using RISP are $0.66$ and $0.94$. We performed a comprehensive analysis of the peak signal-to-noise ratio (PSNR) of the reconstruction results before and after using RISP to optimize (All three data sets are normalized) and using bm4d. The results are as follows. It further demonstrates the effectiveness of the RISP optimization.
@@
Table 1: The PSNR of different results
UBP RISP bm4d
PSNR 4.6875dB 23.3810dB 17.0552dB
In conclusion, in the simulation experiments, the reconstruction results obtained using the UBP algorithm and further optimized by the RISP algorithm demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
Parameters Selection and Computing Environment
@@
Table 2: 3D PA image reconstruction under 1024 hemispherical system
@@
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor $\times$ 128. In UBP algorithm reconstruction, we use Taichi[23, 24] for GPU acceleration. For the reconstruction with PA signal size of 50$\times$4096 and region size of 128$\times$320$\times$256, the time consuming is about 0.5s. The computation of the probability matrix takes about 35 seconds, and the entire iteration process takes about 10 seconds. So the entire RISP optimization process took less than a minute and 10 seconds ($0.5\times 50+35+10$). Of course, we have skills in the choice of parameters, and the choice of $x$ is generally about one-tenth of the total number of elements. We also found that when the initial value of lr is about 0.001, the iterative result can be quickly optimized. And the number of iterations is much lower than $num\_iters$, and about 50.
3.1.2 2D PA Image Reconstruction under 256 Circle System
We then used fundus blood vessels[25] as target to demonstrate the high performance of the RISP algorithm (Figure 4(a)). At a grid point spacing of 0.05 mm, the total grid size of the imaging area is 512$\times$512 (25.6 mm $\times$25.6 mm). The radius of the circle system is 5cm. Then we obtained PA signals collected by these 256 arrays through forward propagation of the K-Wave toolbox.
Comparison of Results before and after Optimization
From the Figure 4, the artifacts significantly decrease after optimization by the RISP algorithm. Details of blood vessels are displayed more clearly. The PSNR results are shown in Table 3. In conclusion, in this experiment, the reconstruction results obtained using the UBP algorithm and further optimized by the RISP algorithm demonstrated excellent performance. Both the artifacts around the internal small blood vessels and the overall peripheral artifacts are weakened. And the details of the results optimized by the bm3d algorithm almost disappear.
Table 3: The PSNR of different results
UBP RISP bm3d
PSNR 12.9899dB 13.4995dB 10.9612dB
Figure 4: Comparison of 2D photoacoustic reconstruction results under 256 circle system. (a) Ground truth. (b) Reconstruction result of UBP algorithm. (c) Reconstruction result after optimization of RISP algorithm.
Parameters Selection and Computing Environment
@@
Table 4: 2D PA image reconstruction under 256 circle system
@@
$num\_iters$ 50
In this experiment, the CPU type is 13th Gen Intel(R)Core(TM)i5-13500H. In UBP algorithm reconstruction, for the reconstruction with PA signal size of 50$\times$2048 and region size of 768$\times$768, the time consuming is about 0.50s. The computation of the probability matrix takes about 30 seconds, and the entire iteration process takes about 10 seconds. So, the entire RISP optimization process took about 3 minutes ($0.50\times 100+30+10$).
@@ 3.2 In Vivo Human Experiment Results
Then in vivo human study data, including human arm was acquired by Li’s work[21] using synthetic matrix array. The system employed a nonfocusing linear array (customized by Imasonics, France) to receive PA signals. The linear array has 256 elements with a pitch of 0.5 mm and a kerf of 0.1 mm, i.e., a total length of 12.8 cm. The center frequency of the ultrasonic array is 3.5 MHz with over 80% bandwidth. PA signal is amplified 1500 times via self-built two-stage amplifier, and then received by a data acquisition system (Marsonics DAQ, Tianjin Langyuan Technology Co., Ltd. China) at 40 MHz sampling rate.
In the original work, a linear array consisting of 256 elements was moved 2969 steps, with a step size of 0.1 mm for each movement, resulting in a large-scale synthetic matrix of 256 × 2969 elements. To create a sparse-view setup for better evaluation of the RISP algorithm’s performance, we extracted 120 rows at equal intervals (approximately 25% of the original data), which corresponds to a step size of 2.5 mm for each scan of the linear array. Therefore, a planar matrix array with $m\times n$ elements was used.
@@ 3.2.2 Comparison of Results before and after Optimization
In this experiment, we utilized the RISP algorithm to optimize the reconstruction results obtained from the UBP algorithm. The detailed comparison of results before and after optimization is presented in Figure 5.
We compared the hand imaging results before and after the RISP optimization of the UBP reconstruction. From the top-view perspective, the optimization significantly reduced noise, and both thick and thin blood vessels were displayed more clearly. From the front-view perspective, in addition to making the vascular information more distinct, the lateral noise at the bottom was also effectively suppressed. And the contrast-to-noise ratio (CNR) using UBP is 8.7648, while the result after using RISP is 53.8096. These findings demonstrate that the reconstruction quality after RISP optimization is exceptionally high.
In conclusion, in vivo human experiment, the reconstruction results obtained using the UBP algorithm and further optimized by the RISP algorithm also demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
@@
(b) Hand’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RISP results.
@@ 3.2.3 Parameters Selection and Computing Environment
Table 5: 3D PA image reconstruction under synthetic matrix array
@@
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor $\times$ 128. In UBP algorithm reconstruction, we use Taichi for GPU acceleration. For the reconstruction with PA signal size of 30720$\times$2048 and region size of 625$\times$600$\times$250, the time consuming is about 10.2s. For the reconstruction with PA signal size of 4000$\times$2048 and region size of 625$\times$600$\times$250, the time consuming is about 3.5s. The computation of the probability matrix takes about 140 seconds, and the entire iteration process takes about 90 seconds. So the entire RISP optimization process took less than 7 minutes ($3.5\times 50+140+90$).
@@ 4 Discussion and Future
In this study, we focus on demonstrating the optimization effectiveness of the RISP algorithm applied to the reconstruction results of the UBP algorithm in the context of 3D imaging systems. Through a series of specific experiments, the RISP algorithm has shown exceptional performance in both simulated studies and live small animal experiments. Notably, from a theoretical perspective, the RISP algorithm is not constrained to a specific imaging dimension, making it applicable to both 3D and 2D imaging scenarios. Moreover, the core concept of the RISP algorithm lies in optimizing existing imaging results rather than directly reconstructing images within the signal domain. This unique characteristic grants the algorithm broad applicability, enabling effective suppression of noise and artifacts caused by transducer-related factors (e.g., transducer distribution and other system properties). Importantly, the RISP optimization is not tied to a specific reconstruction algorithm; other algorithms can also benefit from its enhancement capabilities.
Furthermore, the potential applications of the RISP algorithm extend beyond photoacoustic computed tomography. Theoretical analysis suggests that the algorithm could be generalized to other imaging fields, such as ultrasound imaging, computed tomography (CT), and various other medical imaging or industrial inspection scenarios. In these fields, the sampling and imaging processes of various devices may also introduce artifacts and noise that vary with the properties of the transducers. In such cases, the RISP algorithm is expected to offer significant assistance.
Going forward, we plan to expand our research in two key directions. Firstly, we aim to further investigate the optimization capabilities of the RISP algorithm across a broader spectrum of existing reconstruction algorithms, exploring its adaptability under diverse imaging conditions, modalities, and algorithmic outputs. Secondly, we intend to extend the application of the RISP algorithm to other areas, such as ultrasound imaging, CT, and potentially other advanced medical imaging techniques like magnetic resonance imaging (MRI) and positron emission tomography (PET). These efforts would further validate its extensive practical value and generalizability.
In the long term, we aspire to make the RISP algorithm an essential tool for a wide range of imaging systems. Through continued optimization and extension studies, this algorithm has the potential to provide comprehensive solutions for improving imaging quality, resolution, and reliability across diverse application scenarios.